I need help on solving systems using substitutions.
Y=X+4
Y=3X

Answer by jim_thompson5910(35256)

y=x+4

3x=x+4

3x-x=4

2x=4

x=4/2

x=2

-------------------

y=x+4

y=2+4

y=6


So the solutions are x = 2 and y = 6
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------